OpenStudy (anonymous):

The temperature fell from 0 degrees to 6 3/5 degrees below 0 in 2 1/5 hours. What was the temperature change per hour?

OpenStudy (anonymous):

no

OpenStudy (skullpatrol):

$$\Huge 6\frac{3}{5} \div 2\frac{1}{5}$$

OpenStudy (skullpatrol):

$$\Huge \frac{33}{5} \cdot \frac{5}{11}=3$$

OpenStudy (skullpatrol):

Answer: -3 degrees per hour

OpenStudy (anonymous):

thank you

OpenStudy (skullpatrol):

thanks for asking :-)